### v1

# Mitigating Feature Gap for Adversarial Robustness by Feature Disentanglement

Abstract
Deep neural networks are vulnerable to adversarial samples. Adversarial fine-tuning methods aim to enhance adversarial robustness through fine-tuning the naturally pre-trained model in an adversarial training manner. However, we identify that some latent features of adversarial samples are confused by adversarial perturbation and lead to an unexpectedly increasing gap between features in the last hidden layer of natural and adversarial samples. To address this issue, we propose a disentanglement-based approach to explicitly model and further remove the latent features that cause the feature gap. Specifically, we introduce a feature disentangler to separate out the latent features from the features of the adversarial samples, thereby boosting robustness by eliminating the latent features. Besides, we align features in the pre-trained model with features of adversarial samples in the fine-tuned model, to further benefit from the features from natural samples without confusion. Empirical evaluations on three benchmark datasets demonstrate that our approach surpasses existing adversarial fine-tuning methods and adversarial training baselines.
Mitigating Feature Gap, Feature Disentanglement, Adversarial Robustness, Adversarial Fine-Tuning
Figure 1: Toy illustration of Disentanglement. We model features of adversarial samples and then remove the latent features confused by adversarial perturbation to correct the prediction.
1 Introduction
Deep neural networks (DNNs) have shown impressive performances in various domains of machine learning. However, it has been demonstrated that DNNs are susceptible to adversarial samples and the prediction can be easily manipulated (Goodfellow et al., 2014). Adversarial samples deceive DNNs by introducing imperceptible noise to clean samples. With the widespread use of DNNs, the presence of adversarial samples poses a growing potential threat, making it crucial to enhance the robustness of networks.
Many defensive techniques are proposed including adversarial training (AT) (Madry et al., 2017; Zhang et al., 2019; Wang et al., 2020; Wu et al., 2020; Huang et al., 2020; Dong et al., 2021; Jia et al., 2022; Jin et al., 2022; Zhou et al., 2022), adversarial fine-tuning (AFT) (Suzuki et al., 2023; Maini et al., 2020; Madaan et al., 2021; Croce & Hein, 2022; Jeddi et al., 2020; Zhu et al., 2023; Si et al., 2020; Agarwal et al., 2023; Liu et al., 2023), adversarial purification (Yoon et al., 2021; Sun et al., 2023; Lee & Kim, 2023; Allen-Zhu & Li, 2022; Nie et al., 2022), adversarial detection (Hickling et al., 2023; Zheng & Hong, 2018; Pang et al., 2018), etc. It is universally acknowledged that AT is the most effective defensive technique, but it suffers from several times the training cost as standard (natural) training. To save the cost of training time, AFT employs the same loss function as AT to fine-tune the pre-trained model within a few epochs.
We identify an issue of the increasing feature gap by analyzing three models fine-tuned by three AFT methods (vanilla AFT, AFT+RGKD, ARREST) (Suzuki et al., 2023; Jeddi et al., 2020). As illustrated in Figure 2, the feature distance between natural and adversarial samples abnormally maintains an increasing direction during the fine-tuning process. Though ARREST initially exhibits notable fluctuations with a decreasing trend, the overall magnitudes of feature distance in three AFT methods show an increasing trend. We suppose that some latent features are easy to be confused by adversarial attacks and lead to the growing gap between the features of natural and adversarial samples. Considering that a robust model should treat natural and adversarial samples equally and extract similar features from them, we expect to remove the latent features causing the feature gap to achieve better robust performance.
In this paper, we analyze the latent features causing the feature gap and introduce an approach based on disentanglement to explicitly model and remove these features. We propose a new technique named Adversarial Fine-tuning via Disentanglement (AFD). First, we disentangle the features of adversarial samples as two components, in which the latent features causing the feature gap are identified as the features confused by adversarial perturbation. Next, we propose a feature disentangler to model and separate them from the features of adversarial samples. We maximize the predicted probabilities of the wrong predicted classes to acquire the latent features. Then we impose a constraint to distance the features of adversarial samples from them, to eliminate the latent features and mitigate the feature gap. Besides, we align features of adversarial samples in the fine-tuned model with features of natural samples in the naturally pre-trained model, further eliminating the latent features. Experiments demonstrate that our AFD alleviates the gap between features of natural and adversarial samples and improves the robustness. The key contributions are summarized as follows:
  • We observe the gap in features between natural and adversarial samples anomalously increases when applying adversarial fine-tuning methods, resulting from the latent features confused by adversarial perturbation.
  • We propose a disentanglement-based approach, which explicitly models and further eliminates the latent features causing the feature gap by a disentangler and loss function of alignment.
  • Empirical evaluations show our approach mitigates the gap in features between natural and adversarial samples and surpasses existing methods. We also provide extended analyses for a holistic understanding.
2 Related Work
Adversarial Training Adversarial training (AT) is the most effective technique to defend against attacks. Numerous studies have been conducted to boost robustness and improve the trade-off between accuracy and robustness. Madry et al. (2017) propose Projected Gradient Descent (PGD) attack and PGD-based adversarial training (PGD-AT), compelling the model to correctly classify adversarial samples within the epsilon sphere during training, which is the pioneer in adversarial learning. Zhang et al. (2019) reduce the divergence of probability distributions of natural and adversarial samples to mitigate the difference between robust and natural accuracy. Wang et al. (2020) find that misclassified samples harm adversarial robustness significantly, and propose to improve the model’s attention to misclassification by adaptive weights. Huang et al. (2020) replace labels with soft labels predicted by the model and adaptively reduce the weights of misclassified samples to alleviate the robust overfitting problem. Dong et al. (2021) also propose a similar idea of softening labels and explain the different effects of hard and soft labels on robustness by investigating the memory behavior of the model for random noisy labels. Zhou et al. (2022) embed a label transition matrix into models to infer natural labels from adversarial noise. Though AT has good robust performances, it suffers from a large quantity of computing expenses and training costs. Some researchers have developed approaches to accelerate adversarial training, commonly known as fast adversarial training (FAT) (Wong et al., 2020; Kim et al., 2020; Huang et al., 2023). Since these FAT methods only require a little training time like AFT, we compare them with our method.
Adversarial Fine-tuning Adversarial fine-tuning (AFT) has been employed to enhance the pre-trained model within a few epochs for various targets, particularly for achieving adversarial robustness. Moosavi-Dezfooli et al. (2018) analyze the effect of AFT by comparing the decision boundary of a DNN before and after applying AFT. Some researchers (Tramer & Boneh, 2019; Maini et al., 2020; Madaan et al., 2021; Croce & Hein, 2022) aim to achieve multiple $l_{p}$ adversarial robustness through specific strategies. Jeddi et al. (2020) employ vanilla AFT to expedite the training of robust models with a few training epochs and a warm-up learning rate schedule. Based on representation learning, Suzuki et al. (2023) propose ARREST and AFT+RGKD, which reduce the feature gap between natural and adversarial samples. Zhu et al. (2023) propose a metric to measure the robustness of modules and fine-tune them for better out-of-distribution performance. Liu et al. (2023) introduce a two-pipeline structure to improve the relationship between the weight norm and its gradient norm in batch normalization layers. Despite the extensive work, few methods (vanilla AFT, ARREST, and AFT+RGKD) are based on naturally pre-trained models, which are important baselines in our paper.
3 Methodology
In this section, we first reveal the issue of the increasing gap of features between natural and adversarial samples in AFT. Then we introduce a disentanglement-based method, to model and further remove the latent features leading to the feature gap. Our method is expected to mitigate the feature gap to enhance adversarial robustness.
3.1 Preliminaries
We denote a DNN as $h(x)=\omega\cdot g(x;\theta_{g})$, where $x$ is the input image, and $\theta_{g}$ is the parameters of the feature extractor $g$, i.e., the model before the last linear classifier. $\omega$ denotes the parameters of the last linear classifier. The parameters of the naturally pre-trained and fine-tuned model are denoted as $\theta^{P}$ and $\theta^{F}$, respectively. The architecture and dataset for both pre-training and fine-tuning remain consistent, and the model parameters $\theta^{F}$ are initialized by $\theta^{P}$. $x$ denotes the clean sample, and $x^{\prime}$ denotes the adversarial sample. ${\|{\delta}\|}_{p}\leq\epsilon$ denotes the adversarial perturbation, where $\epsilon$ represents the budget of adversarial attack. We use ${\|{\cdot}\|}_{p}$ to denote the $l_{p}$ norm. Adversarial and natural features denote the features in the last hidden layer of adversarial and natural samples, i.e., $g(x^{\prime};\theta_{g})$ and $g(x;\theta_{g})$ respectively. $\mathcal{L}$ denotes the cross-entropy loss.
3.2 Model the Latent Feature causing Feature Gap
We evaluate the feature gap in AFT by measuring the ${L}_{p}$ distance between natural and adversarial features. As shown in Figure 2, the overall magnitudes of feature distance between natural and adversarial samples in three AFT methods (vanilla AFT, ARREST, AFT+RGKD) are becoming larger. The gap between the two types of features in ARREST initially decreases and then increases, which exhibits a V-shape trend. In addition, the feature gap in other AFT methods presents a consistently increasing trend. Despite AFT enhancing robust accuracy during training, the feature distance in AFT maintains a growing trend overall. Intuitively, the ideal robust model should have a consistent understanding and analysis of natural and adversarial samples, thus extracting similar (the same) features from them. Figure 2 indicates that existing AFT methods neglect the gap of features between natural and adversarial samples. Narrowing the feature gap is expected to achieve better performance and enhance robust accuracy.
To realize the target, we can disentangle two types of features from the features of natural and adversarial samples. We define natural features $f_{nat}$ as the unconfused features $f_{u}$. We can define adversarial features $f_{adv}$ as the combination of the unconfused features $f_{u}$ and the confused features $f_{c}$, the latter of which leads to the gap in features between natural and adversarial samples. In this modeling, the unconfused features represent the features that contribute to accurate classification, which are easily extracted from natural samples. Besides, the confused features are identified as the latent features causing the gap between features of natural and adversarial samples, which are confused by the adversarial perturbation. As illustrated in Figure 1, if the confused features are removed in adversarial features, the features of adversarial samples can align with those of natural samples, resulting in a zero feature gap and accurate predictions for adversarial samples.
Figure 2: $L_{p}$ distances of features between natural and adversarial samples on ResNet18 on CIFAR-10. The horizontal axis denotes the training epoch. Despite the initial decreasing fluctuations observed in ARREST and AFT+RGKD, AFT methods exhibit an increasing trend on the whole.
We suppose the phenomenon in Figure 2 is attributable to the presence of the confused features $f_{c}$. Because the confused features persist as latent features and don’t diminish during fine-tuning, the gap between features of natural and adversarial samples cannot be bridged. The confused features induce the gap between natural and adversarial features, leading to the limited robust performance. Based on our observations and analyses, we introduce a new approach named Adversarial Fine-tuning via Disentanglement (AFD), as shown in Figure 3. It applies a feature disentangler to explicitly extract the confused features, and remove them by the elimination and alignment of features.
Figure 3: Overview of our AFD. Lock denotes the frozen parameters. The disentangler disentangles adversarial features into $f_{1}$ and $f_{2}$. We obtain confused features $f_{2}$ by matching the wrong predicted labels in $\mathcal{L}_{1}$. Next, we employ $\mathcal{L}_{2}$ to eliminate confused features from unconfused features $f_{1}$. Besides, we apply $\mathcal{L}_{3}$ to further align $f_{1}$ with pre-trained unconfused features $f_{nat}^{\prime}$ to correct the prediction.
3.3 Adversarial Fine-tuning Based on Disentanglement
The primary strategy is to explicitly disentangle the unconfused and confused features from the features of adversarial samples. We propose a feature disentangler to model confused features for further elimination. The features in the last hidden layer are input into the disentangler, yielding two disentangled feature vectors. We define these feature vectors as 1) $f_{1}$, features contributing to correct classification without confusion; and 2) $f_{2}$, features inducing wrong predictions of adversarial samples. To implement this disentangling, we introduce a disentangler with two linear blocks $D_{1},D_{2}$ between the feature extractor $g$ and the linear classifier $\omega$. These linear blocks are expected to disentangle adversarial features into the unconfused features and confused features. The disentangled feature vectors $f_{1}$ and $f_{2}$ are constrained to match the unconfused features and confused features, formulated as:
$\begin{split}f_{1}&=\ {D_{1}}(g(x^{\prime};\theta_{g}^{F});\theta_{D_{1}}),\\ f_{2}&=\ {D_{2}}(g(x^{\prime};\theta_{g}^{F});\theta_{D_{2}}),\\ {\rm where}\ &\left\{\begin{array}[]{lr}f_{1}\stackrel{{\scriptstyle D_{1}}}{{\longrightarrow}}f_{u},&\\ f_{2}\stackrel{{\scriptstyle D_{2}}}{{\longrightarrow}}f_{c},&\end{array}\right.\end{split}$ (1)
where $\theta_{g}^{F}$ denotes the parameters of the feature extractor in the fine-tuned model. We take $f_{1}$ to correctly classify samples and $f_{2}$ to match the confused features. The input and output dimensions of the linear block $D_{1},D_{2}$ are set to match the channel dimensions of the feature vector of adversarial samples before disentanglement. This ensures the parameters of the linear block $\theta_{D_{1}}$ can be seamlessly integrated with the parameter of the last linear classifier $\omega^{F}$, expressed as $\theta_{D_{1}}\cdot\omega^{F}={\omega^{F\prime}}$. Thus we only need to load the linear classifier $\omega^{F\prime}$ without additional disentangling modules in the deployment (test) phase.
To achieve good performance on both accuracy and robustness, we should design suitable loss functions to realize Formula 1. First, we suggest fitting $f_{2}$ to incorrectly predicted labels $y^{\prime}$ to model the confused features. The confused features result in obvious shifts in both features and predicted probabilities, leading to wrong predicted class labels in most cases. Thus, it can be identified as the latent features causing the gap between natural and adversarial features. To disentangle the confused features from adversarial features, we force the disentangled features $f_{2}$ to match incorrect class labels, formulated as:
$\begin{split}\mathcal{L}_{1}=\ &{\rm\mathop{arg\ min}\limits_{\theta_{D_{2}}}}\ \mathcal{L}(\omega^{F}\cdot f_{2},y^{\prime}),\\ {\rm where}\quad&y^{\prime}={\rm\mathop{arg\ sort}\limits_{y^{\prime}\neq y}}\ \omega^{F}\cdot f_{1},\end{split}$ (2)
where $y^{\prime}$ denotes wrong predicted labels with the maximum predicted probability, and sort represents finding the index of the maximum value. We only optimize the disentangling block $D_{2}$ instead of the entire model. This selective optimization prevents potential harm to the feature extractor that could result from learning the confused features.
In addition, we introduce a constraint on $f_{1}$ to keep away from the confused features $f_{2}$, removing the confused features from the features of adversarial samples. The exclusive force between $f_{1}$ and $f_{2}$ encourages the elimination of the confused features from the features of adversarial samples. This constraint aims to protect adversarial features from being perturbed by the confused features, and then gain the unconfused features $f_{1}$. Given by:
$\mathcal{L}_{2}={\rm\mathop{arg\ max}\limits_{\theta^{F},\theta_{D_{1}}}}\ \mathcal{D}(f_{1},\bar{f_{2}}),$ (3)
where $\mathcal{D}$ denotes a distance function, $\bar{f_{2}}$ means the gradient back-propagation of $f_{2}$ is frozen.
3.4 Adversarial Fine-tuning Based on Alignment
The second strategy is to align the adversarial features in the fine-tuned model with the natural features in the pre-trained model. Since the pre-trained model has a high natural accuracy and doesn’t suffer from adversarial attacks, the natural features in the pre-trained model can be identified as ideal unconfused features. Thus, adversarial features in the fine-tuned model can be greatly improved by being aligned with the natural features in the pre-trained model. Besides, lots of work (Zhang et al., 2019; Wang et al., 2020; Suzuki et al., 2023) conducts an alignment between natural and adversarial features, which is regarded to mitigate the trade-off between accuracy and robustness. We take the disentangled features $f_{1}$ to match the natural features in the pre-trained model, further removing the confused features from $f_{1}$. The AFT loss function with the alignment can be formulated as:
$\mathcal{L}_{3}={\rm\mathop{arg\ min}\limits_{\theta^{F},\theta_{D_{1}}}}\ \mathcal{L}(\omega^{F}\cdot f_{1},y)+\gamma\cdot\mathcal{D}(f_{1},g(x;\theta_{g}^{P})),$ (4)
where $\theta_{g}^{P}$ denotes parameters of the feature extractor in the pre-trained model, and $\gamma$ denotes the weight of the alignment. Therefore, the objective function of our AFD is defined as follows:
$\mathcal{L}_{total}=\alpha\cdot\mathcal{L}_{1}+\beta\cdot\mathcal{L}_{2}+\mathcal{L}_{3},$ (5)
where $\alpha$ and $\beta$ denote the weights of $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$. The algorithm of AFD is as follows:
Algorithm 1 Adversarial Fine-tuning via Disentanglement
0:  Training set $\mathcal{D}$, The $\mathcal{L}_{1}$ weight $\alpha$, The $\mathcal{L}_{2}$ weight $\beta$, Number of epochs $T$, Learning rate $\eta$, Inner maximization $\mathcal{A}$, Naturally pre-trained model parameters $\theta^{P}$, Model parameters except linear block $D_{2}$ parameters $\theta_{1}$, Linear block $D_{2}$ parameters $\theta_{2}$.
0:  Fine-tuned model parameters $\theta_{1}$.
  Initialize $\theta_{1}$ by $\theta^{P}$;
  for $t=1,\cdots,T$ do
     for $x,y\in\mathcal{D}$ do
        $x^{\prime}=\mathcal{A}(x)$;
        Calculate features $f_{1},f_{2}$ by Formula 1;
        Calculate loss $\mathcal{L}_{1}$ by Formula 2;
        $\theta_{2}=\theta_{2}-\eta\cdot\frac{\mathrm{d}(\alpha\cdot\mathcal{L}_{1})}{\mathrm{d}\theta_{2}}$;
        Calculate loss $\mathcal{L}_{2}$ by Formula 3;
        $\theta_{1}=\theta_{1}-\eta\cdot\frac{\mathrm{d}(\beta\cdot\mathcal{L}_{2})}{\mathrm{d}\theta_{1}}$;
        Calculate loss $\mathcal{L}_{3}$ by Formula 4;
        $\theta_{1}=\theta_{1}-\eta\cdot\frac{\mathrm{d}\mathcal{L}_{3}}{\mathrm{d}\theta_{1}}$;
     end for
  end for
4 Experiments
This section presents experiments with AFD. We first introduce our experiment setting. We further make comparisons on multiple architectures and datasets to show our superiority. Then we conduct an ablation study to show the functions of the disentangling module and alignment. Finally, we conduct an empirical analysis to support our hypothesis.
4.1 Setting
Datasets We conduct our experiments on three benchmark datasets including on CIFAR-10 and CIFAR-100 (Krizhevsky, 2009), Tiny-ImageNet (Deng et al., 2009). CIFAR-10 dataset contains 60,000 color images having a size of $32\times 32$ in 10 classes, with 50,000 training and 10,000 test images. CIFAR-100 dataset contains 50,000 training and 10,000 test images in 100 classes. Tiny-ImageNet dataset contains 100000 images of 200 classes (500 for each class) downsized to 64×64 colored images. Each class has 500 training images, 50 validation images and 50 test images. We only use training and validation images of Tiny-ImageNet in experiments.
Baselines To make a comprehensive comparison, we have employed various techniques to demonstrate our superiority. Our primary baselines are AFT methods including vanilla AFT (vAFT) (Jeddi et al., 2020), AFT+RGKD (AFKD) (Suzuki et al., 2023), and ARREST (Suzuki et al., 2023). All of them are built upon naturally pre-trained models. Additionally, we compare our AFD with state-of-the-art AFT methods such as RiFT (Zhu et al., 2023) and TWINS (Liu et al., 2023), which rely on adversarially pre-trained models. Besides, we select several advanced fast adversarial training methods for comparison, including FreeAT (Shafahi et al., 2019), FGSM-GA (Andriushchenko & Flammarion, 2020), ATAS (Huang et al., 2023), since they are also low-cost techniques. Moreover, we compare with baseline methods of adversarial training including PGD-AT (Madry et al., 2017), TRADES (Zhang et al., 2019) to show our great advance in robustness.
Optimization Details The optimizer used in all the experiments is Stochastic Gradient Descent (SGD) optimizer with a momentum of 0.9. We naturally pre-train all the models with a batch size of 128, training epochs of 100, and a weight decay of ${5.0\times 10^{-4}}$. The learning rate starts at 0.1 and then decays by $\times$0.1 with transition epochs $\{75,90\}$, following Zhang et al. (2019). For AT methods, we train all the models with a batch size of 128, training epochs of 120, and a weight decay of ${2.0\times 10^{-4}}$. The learning rate starts at 0.1 and then decays by $\times$0.1 with transition epochs $\{75,90,100\}$, following Wang et al. (2020). For our AFD, we fine-tune the pre-trained models with a batch size of 128, an initial learning rate of 0.0025, a weight decay of ${5.0\times 10^{-4}}$, and training epochs of 20. The learning rate is changed to 0.0020 at the 11th epoch and is divided by half every two epochs thereafter, as Suzuki et al. (2023).
[…]
4.2 Main Results against White-Box Attack
To evaluate AFD’s effectiveness, we conducted comparative experiments with other methods on CIFAR-10, CIFAR-100, and Tiny-ImageNet. We also make additional experiments on ResNet50 to fairly compare with TWINS.
CIFAR-10 and CIFAR-100 As shown in Tables 1, methods above the double line denote AT and FAT methods, and methods under the double line denote AFT methods. It shows that among AFT methods, AFD exhibits the best robust performance and ARREST has the second best robust performance on ResNet18. Robust accuracies of AFD against AutoAttack (AA) surpass those of ARREST by 1.27% and 2.58% on CIFAR-10 and CIFAR-100, respectively. As indicated in Table 2, AFD exhibits the best robust performance and AFKT has the second best robust performance among all the methods on WideResNet-34-10. AFD achieves superior robust accuracies against AA compared to AFKD, with improvements of 2.05% and 2.34% on CIFAR-10 and CIFAR-100, respectively. In particular, RiFT doesn’t show great robust performances, since it is fine-tuned with natural data to enhance out-of-distribution performance.
Besides, AFD with fewer training epochs outperforms AT methods. AFD surpasses PGD-AT in all respects. TRADES achieves a larger robust accuracy against AA on CIFAR-10 on ResNet18 than AFD by 1.02%. However, the situation is reversed on CIFAR-100 or on WideResNet-34-10, where AFD outperform TRADES against AA by 0.43%, 2.40%, and 2.37%. This can be attribute to the superiority of AFD and the robust over-fitting in AT. Moreover, FAT methods distinctly lag behind AFD in both natural and robust performances. Despite the advanced robust performance of AFD, its natural accuracy doesn’t surpass state-of-the-art AFT methods (ARREST). We make further discussion in the Section 4.3. In summary, AFD demonstrates superb robustness compared to AT, FAT and AFT methods.
Table 1: Quantitative evaluations of AFD and other methods for CIFAR-10 and CIFAR-100 on ResNet18.
Methods CIFAR-10 CIFAR-100
Clean PGD AA Clean PGD AA
PGD-AT 84.43 48.82 43.68 59.09 24.05 20.67
TRADES 82.37 53.27 48.52 55.47 28.12 23.49
${}^{{\dagger}}$ 78.37 40.90 36.00 50.56 19.57 15.09
${}^{{\dagger}}$ 80.10 49.14 43.44 50.61 24.48 19.42
${}^{{\dagger}}$ 81.22 50.03 45.38 55.49 27.68 22.62
vAFT 83.93 51.40 45.48 60.83 26.55 21.96
RiFT 84.49 49.24 43.66 59.51 24.01 20.48
AFKD 84.71 51.32 46.09 65.54 28.09 22.92
ARREST 85.71 51.23 46.23 67.00 26.89 21.34
AFD 84.66 52.63 47.50 62.94 29.27 23.92
Table 2: The performance of AFD and other methods on CIFAR-10 and CIFAR-100 on WideResNet-34-10.
Methods CIFAR-10 CIFAR-100
Clean PGD AA Clean PGD AA
PGD-AT 86.35 49.57 45.81 60.30 25.14 22.65
TRADES 85.68 52.53 49.22 57.67 28.26 25.25
vAFT 87.42 53.57 48.49 64.92 27.19 23.59
RiFT 84.49 49.24 45.83 60.59 25.57 22.67
AFKD 88.54 54.86 49.57 68.50 30.12 25.28
ARREST 88.42 54.16 49.45 69.52 29.47 24.55
AFD 88.08 56.12 51.62 68.51 32.63 27.62
Figure 4: The clean accuracy (NAT) and robust accuracy against PGD-20 (PGD) of AFD with different hyperparameters.
Tiny-ImageNet We further conduct comparative experiments on a larger dataset, Tiny-ImageNet. As shown in Table 3, on ResNet18, AFD consistently outperforms all the methods in terms of robust performance and the robust performance of TRADES ranks second. The robust accuracy of AFD against PGD exceeds that of TRADES by 1.59% and its robust accuracy against AA outpaces that of TRADES by 1.58%. Moreover, on WideResNet-34-10, AFD keeps the best robust performance and AFKD exhibits the second best robust performance. The robust accuracy of AFD against PGD on WideResNet-34-10 surpasses that of AFKD by 2.56% and its robust accuracy against AA on WideResNet-34-10 is higher than that of AFKD by 2.76%. These findings indicate the superiority of AFD in terms of adversarial robustness on large datasets.
Table 3: The performances of AFD and other methods on Tiny-ImageNet with ResNet18 and WideResNet-34-10 models.
Methods ResNet18 WideResNet-34-10
Clean PGD AA Clean PGD AA
PGD-AT 52.12 19.38 16.79 54.24 19.03 16.51
TRADES 49.28 22.59 17.06 49.22 23.33 18.51
vAFT 52.75 21.29 16.76 57.17 22.24 18.02
RiFT 52.24 20.89 17.05 55.09 19.51 16.79
AFKD 57.17 22.29 16.68 60.92 24.27 19.01
ARREST 60.36 19.28 12.58 65.07 22.73 16.67
AFD 55.58 24.18 18.64 59.58 26.83 21.77
ResNet50 We follow Liu et al. (2023) to make comparisons on ResNet50. Notice that TWINS employs a model adversarially pre-trained on ImageNet (Deng et al., 2009), and the number of its training periods is 60. We set the training epochs of AFD as 40. Results are shown in Table 4. Although TWINS benefits from prior knowledge of ImageNet and adversarial robustness of the pre-trained model, AFD still surpasses its robust performance on both CIFAR-10 and CIFAR-100. AFD achieves improvements of 1.54% and 1.56% under AA compared to TWINS, and its robust accuracies against PGD surpass those of TWINS by 1.88% and 2.52%. This demonstrates its superiority in terms of adversarial robustness on ResNet50.
Table 4: The performances of AFD and the state-of-the-art method TWINS on CIFAR-10 and CIFAR-100 on ResNet50.
Methods CIFAR-10 CIFAR-100
Clean PGD AA Clean PGD AA
${}^{{\dagger}}$ 89.77 52.24 48.46 69.48 28.52 23.47
${}^{{\dagger}}$ 91.95 52.46 49.02 72.12 29.12 25.72
AFD 86.31 54.34 50.56 65.48 31.64 27.28
In summary, these experiments indicate that AFD always outperforms all the baselines for different adversarial attacks, datasets, and network architectures in the aspect of robustness.
4.3 Ablation Study
Effect of Disentangler As mentioned in Section 3.3, the disentangler should disentangles adversarial features into the confused and unconfused features, and further eliminate the confused features. As shown in Table 5, vAFT with the disentangler (vAFT+D) exhibits superior robustness compared to vAFT by 1.28% against PGD and by 1.40% against AA, respectively. It shows the feature disentangler effectively eliminates the confused features that lead to wrong predictions. Moreover, the natural accuracy of AFT-D outperforms that of AFT by 0.70%. This could be attributed to the disentangling module removing the class-irrelevant features of natural samples such as background and shadow.
Effect of Alignment As shown in Table 5, vAFT with the alignment of features (vAFT+A) exhibits superior robustness compared to AFT against both PGD and AA by 0.39% and 0.72%, with a marginal increase in natural accuracy by 0.67%. It indicates that the alignment successfully assists to further remove the confused features.
Effect of Combination The combination of disentangler and alignment (AFD) has the best performance on both natural and robust accuracy, which shows two strategies are compatible and the combination can boost the positive effects of individual strategies. However, in Table 1 and other tables, AFD exhibits a lower natural accuracy than state-of-the-art AFT methods. This is because the unconfused features $f_{1}$ have not been trained with natural samples, which may lead to an accuracy-robust trade-off.
Table 5: The performances of vAFT and other methods on CIFAR-10 on ResNet18. ‘+D’ denotes the method with the disentangler, and ‘+A’ denotes the method with the alignment of features.
Methods vAFT vAFT+D vAFT+A AFD
Clean 83.93 84.63 84.61 84.66
PGD 51.40 52.68 51.71 52.63
AA 45.48 46.88 46.81 47.50
Figure 5: $L_{p}$ distances of features between natural and adversarial samples. The horizontal axis represents the training epoch.
Effect of Hyperparameter There are three hyperparameters in our AFD: $\alpha,\beta$ and $\gamma$. They control the weights of $\mathcal{L}_{1}$, $\mathcal{L}_{2}$, and the distance function in $\mathcal{L}_{3}$, respectively. We have conducted experiments with various combinations of weights on ResNet18 on CIFAR-10, as illustrated in Figure 4. There is a remarkable trade-off between accuracy and robustness when $\beta$ and $\gamma$ are growing. On the one hand, a larger value of $\beta$ indicates a larger exclusive force between the unconfused and confused features, leading to less non-robust features utilized in classification and better robustness. On the other hand, as $\gamma$ increases, the natural accuracy increases but robust accuracy drops. This occurs because more feature information of natural samples is transferred to the fine-tuned model through the feature alignment. However, the variations of both natural and robust accuracy are less than 0.60%, demonstrating that our AFD is insensitive to variations of hyperparameters.
4.4 Empirical Analysis of Disentangled Features
We evaluate the performances of two types of disentangled features $f_{1}$, $f_{2}$ on various datasets and models. As shown in Table 6, the natural accuracies of $f_{2}$ are lower than those of $f_{1}$ by $4.53\%\sim 45.05\%$. It indicates the disentangler has successfully extracted and removed the confused features as $f_{2}$. High accuracies of $f_{1}$ indicates models benefit from eliminating the confused features. Interestingly, the disparity of accuracy on CIFAR-100 is smaller than that on CIFAR-10. We suppose that a dataset with more classes has more complex feature space, and it is more difficult to extract confused features by the linear disentangler. Maybe nonlinear disentanglers work better.
Table 6: The natural accuracies of the unconfused features $f_{1}$ and confused features $f_{2}$. RN and WRN are abbreviations of ResNet and WideResNet, respectively. $\Delta$ denotes the accuracy disparity.
Model Data $f_{1}$ $f_{2}$ $\Delta$
RN18 CIFAR-10 84.66 39.61 45.05
RN18 CIFAR-100 62.94 58.41 4.53
RN50 CIFAR-10 83.18 47.35 35.83
RN50 CIFAR-100 61.68 54.97 6.71
WRN-34-10 CIFAR-10 88.08 46.89 41.19
WRN-34-10 CIFAR-100 68.51 62.84 5.67
Moreover, we analyze the feature distance between the unconfused features $f_{1}$ and natural features $f_{nat}$, as illustrated in Figure 5. In contrast to the growing trend of feature distance in existing AFT methods, the optimization trajectory of the feature distance in AFD exhibits a fast convergence trend. It implies that AFD has mitigated the gap between natural and adversarial features. Additionally, the magnitudes of the feature distance in AFD are smaller than other AFT methods, e.g., the $L_{\infty}$ distance of AFD (0.214) is notably less than those of ARREST and vanilla AFT (0.453, 0.236). These results demonstrates our disentanglement-based approach can enhance the adversarial robustness by alleviating the feature gap.
5 Conclusion
This paper uncovers a surprising increasing trend in the gap of features between natural and adversarial samples in AFT methods, and further investigates it from the perspective of features. We model features as two feature components, in which the confused features are defined as the latent features leading to the feature gap. Then we propose Adversarial Fine-tuning via Disentanglement (AFD) to bridge the feature gap to enhance robustness. Our approach contains two strategies: disentanglement and alignment. We design the feature disentangler to explicitly separate out the confused features from the features of adversarial samples, while the second strategy aligns the adversarial features with the natural features in the pre-trained model. Experiments demonstrate that AFD effectively mitigates the feature gap and achieves significant robust performance, even surpasses adversarial training methods. Future work will aim at the accuracy-robustness trade-off.
Appendix A Performances of Pre-trained models
Table 7: The natural accuracy of the naturally pre-trained models.
Model Data Clean
ResNet18 CIFAR-10 94.60
ResNet18 CIFAR-100 76.55
ResNet18 Tiny-ImageNet 64.81
ResNet50 CIFAR-10 94.30
ResNet50 CIFAR-100 77.20
WideResNet-34-10 CIFAR-10 95.53
WideResNet-34-10 CIFAR-100 80.28
WideResNet-34-10 Tiny-ImageNet 67.86
We conduct pre-training experiments on multiple datasets and architectures. As shown in Table 7, all of the pre-trained models have much higher accuracies than the fine-tuned models. It indicates that it is reasonable for feature alignment to regard features of the natural samples in the pre-trained models as the ideal unconfused features.
Appendix B Feature Visualization
Figure 6: t-SNE visualization of different methods. The proposed AFD shows a clearer separation in the area of the red box.
To further demonstrate the effectiveness of AFD, Figure 6 visualizes the features of AFT methods using t-SNE (Van der Maaten & Hinton, 2008) on ResNet18 on CIFAR-10. We color each point to match its ground-truth label and highlight focused areas with red boxes. It shows the disentangled features learned by AFD have a much clearer class boundary than those learned with other methods. This indicates that AFD successfully mitigates the confusion between features of distinct classes, leading to a more robust prediction.
Appendix C Environments and Time of Experiments
Table 8: The training time (second) of various methods on ResNet18 on CIFAR10. We show the time for one epoch (${\rm Time_{one}}$) and the expected values of the total time (${\rm Time_{total}}$).
Method Epoch ${\rm Time_{one}}$ ${\rm Time_{total}}$
Standard 100 6.0 590
PGD-AT 120 37 4440
${}^{{\dagger}}$ 4428
TARDES 120 53 6360
${}^{{\dagger}}$ 10 119 1188
${}^{{\dagger}}$ 30 68 2052
${}^{{\dagger}}$ 30 36 1080
vanilla AFT 20 37 1330
RiFT 10 25 4690
AFKD 20 40 1390
ARREST 20 36 1310
AFD 20 46 1510
[…]
As shown in Table 8, TRADES costs the most time and ATAS costs the least time among all the methods. The time cost of RiFT surpasses those of other AFT methods because it needs an adversarially pre-trained model. Our AFD costs less time than TRADES by 4850 seconds and its total training time outperforms that of vanilla AFT by 180 seconds. Considering its advanced robustness, the time cost is worthy and acceptable.

### v2

@@
 Abstract
-Deep neural networks are vulnerable to adversarial samples.
-Adversarial fine-tuning methods aim to enhance adversarial robustness through fine-tuning the naturally pre-trained model in an adversarial training manner.
-However, we identify that some latent features of adversarial samples are confused by adversarial perturbation and lead to an unexpectedly increasing gap between features in the last hidden layer of natural and adversarial samples.
-To address this issue, we propose a disentanglement-based approach to explicitly model and further remove the latent features that cause the feature gap.
-Specifically, we introduce a feature disentangler to separate out the latent features from the features of the adversarial samples, thereby boosting robustness by eliminating the latent features.
-Besides, we align features in the pre-trained model with features of adversarial samples in the fine-tuned model, to further benefit from the features from natural samples without confusion.
+Adversarial fine-tuning methods enhance adversarial robustness via fine-tuning the pre-trained model in an adversarial training manner.
+However, we identify that some specific latent features of adversarial samples are confused by adversarial perturbation and lead to an unexpectedly increasing gap between features in the last hidden layer of natural and adversarial samples.
+To address this issue, we propose a disentanglement-based approach to explicitly model and further remove the specific latent features.
+We introduce a feature disentangler to separate out the specific latent features from the features of the adversarial samples, thereby boosting robustness by eliminating the specific latent features.
+Besides, we align clean features in the pre-trained model with features of adversarial samples in the fine-tuned model, to benefit from the intrinsic features of natural samples.
 Empirical evaluations on three benchmark datasets demonstrate that our approach surpasses existing adversarial fine-tuning methods and adversarial training baselines.
-Mitigating Feature Gap, Feature Disentanglement, Adversarial Robustness, Adversarial Fine-Tuning
-Figure 1: Toy illustration of Disentanglement. We model features of adversarial samples and then remove the latent features confused by adversarial perturbation to correct the prediction.
-1 Introduction
+Introduction
 Deep neural networks (DNNs) have shown impressive performances in various domains of machine learning.
-However, it has been demonstrated that DNNs are susceptible to adversarial samples and the prediction can be easily manipulated (Goodfellow et al., 2014).
+However, it has been demonstrated that DNNs are susceptible to adversarial samples and the prediction can be easily manipulated (Goodfellow, Shlens, and Szegedy 2014).
 Adversarial samples deceive DNNs by introducing imperceptible noise to clean samples.
-With the widespread use of DNNs, the presence of adversarial samples poses a growing potential threat, making it crucial to enhance the robustness of networks.
-Many defensive techniques are proposed including adversarial training (AT) (Madry et al., 2017; Zhang et al., 2019; Wang et al., 2020; Wu et al., 2020; Huang et al., 2020; Dong et al., 2021; Jia et al., 2022; Jin et al., 2022; Zhou et al., 2022), adversarial fine-tuning (AFT) (Suzuki et al., 2023; Maini et al., 2020; Madaan et al., 2021; Croce & Hein, 2022; Jeddi et al., 2020; Zhu et al., 2023; Si et al., 2020; Agarwal et al., 2023; Liu et al., 2023), adversarial purification (Yoon et al., 2021; Sun et al., 2023; Lee & Kim, 2023; Allen-Zhu & Li, 2022; Nie et al., 2022), adversarial detection (Hickling et al., 2023; Zheng & Hong, 2018; Pang et al., 2018), etc.
-It is universally acknowledged that AT is the most effective defensive technique, but it suffers from several times the training cost as standard (natural) training.
-To save the cost of training time, AFT employs the same loss function as AT to fine-tune the pre-trained model within a few epochs.
-We identify an issue of the increasing feature gap by analyzing three models fine-tuned by three AFT methods (vanilla AFT, AFT+RGKD, ARREST) (Suzuki et al., 2023; Jeddi et al., 2020).
-As illustrated in Figure 2, the feature distance between natural and adversarial samples abnormally maintains an increasing direction during the fine-tuning process.
+The presence of adversarial samples poses a growing potential threat (Madry et al.
+2017; Croce and Hein 2020; Yu, Gao, and Xu 2021; Wei et al.
+2022, 2023; Liu et al.
+2024; Xia et al.
+2024), making it crucial to enhance the robustness of networks.
+Figure 1: Illustration of our motivation.
+(a) $L_{\infty}$ distances of features between natural and adversarial samples during fine-tuning.
+Previous AFT methods exhibit an increasing trend of the feature gap.
+(b) Toy illustration of Disentanglement.
+We model features of adversarial samples and then remove the specific latent features confused by adversarial perturbation to correct the prediction.
+Many defensive techniques are proposed including adversarial training (AT) (Madry et al.
+2017; Zhang et al.
+2019; Wang et al.
+2020; Wu, Xia, and Wang 2020; Huang, Zhang, and Zhang 2020; Dong et al.
+2021; Jia et al.
+2022; Jin et al.
+2022; Yu et al.
+2022; Zhou et al.
+2022), adversarial fine-tuning (AFT) (Suzuki et al.
+2023; Maini, Wong, and Kolter 2020; Madaan, Shin, and Hwang 2021; Croce and Hein 2022; Jeddi, Shafiee, and Wong 2020; Zhu et al.
+2023; Si et al.
+2020; Agarwal et al.
+2023; Liu et al.
+2023), adversarial purification (Yoon, Hwang, and Lee 2021; Sun et al.
+2023; Lee and Kim 2023; Allen-Zhu and Li 2022; Nie et al.
+2022), adversarial detection (Hickling, Aouf, and Spencer 2023; Zheng and Hong 2018; Pang et al.
+2018), etc.
+AT is regarded as the most effective defensive technique among them, but it suffers from several times the training cost as standard (natural) training.
+To save the cost of training time, AFT employs the same loss function as AT to fine-tune the pre-trained model within a few epochs, which has aroused wide attention.
+We identify an issue of the increasing feature gap by analyzing three models fine-tuned by three AFT methods: vanilla AFT (Jeddi, Shafiee, and Wong 2020), AFT+RGKD, ARREST (Suzuki et al.
+2023).
+As illustrated in Figure 1 (a), the feature distance between natural and adversarial samples abnormally maintains an increasing direction during the fine-tuning process.
 Though ARREST initially exhibits notable fluctuations with a decreasing trend, the overall magnitudes of feature distance in three AFT methods show an increasing trend.
-We suppose that some latent features are easy to be confused by adversarial attacks and lead to the growing gap between the features of natural and adversarial samples.
-Considering that a robust model should treat natural and adversarial samples equally and extract similar features from them, we expect to remove the latent features causing the feature gap to achieve better robust performance.
-In this paper, we analyze the latent features causing the feature gap and introduce an approach based on disentanglement to explicitly model and remove these features.
+We suppose that some specific latent features are easily confused by adversarial attacks and lead to the growing gap between the features of natural and adversarial samples.
+Considering that a robust model should treat natural and adversarial samples equally and extract similar features from them, we expect to remove the specific latent features causing the feature gap to achieve better robust performance, as shown in Figure 1 (b).
+In this paper, we analyze the specific latent features causing the feature gap and introduce an approach based on disentanglement to explicitly model and remove these features.
 We propose a new technique named Adversarial Fine-tuning via Disentanglement (AFD).
-First, we disentangle the features of adversarial samples as two components, in which the latent features causing the feature gap are identified as the features confused by adversarial perturbation.
+First, we disentangle the features of adversarial samples as two components, in which the specific latent features causing the feature gap are identified as the features confused by adversarial perturbation.
 Next, we propose a feature disentangler to model and separate them from the features of adversarial samples.
-We maximize the predicted probabilities of the wrong predicted classes to acquire the latent features.
-Then we impose a constraint to distance the features of adversarial samples from them, to eliminate the latent features and mitigate the feature gap.
-Besides, we align features of adversarial samples in the fine-tuned model with features of natural samples in the naturally pre-trained model, further eliminating the latent features.
-Experiments demonstrate that our AFD alleviates the gap between features of natural and adversarial samples and improves the robustness.
-The key contributions are summarized as follows:
-  • We observe the gap in features between natural and adversarial samples anomalously increases when applying adversarial fine-tuning methods, resulting from the latent features confused by adversarial perturbation.
-  • We propose a disentanglement-based approach, which explicitly models and further eliminates the latent features causing the feature gap by a disentangler and loss function of alignment.
+We maximize the predicted probabilities of the wrong predicted classes to acquire the specific latent features.
+Then we impose a constraint to distance the features of adversarial samples from them, to eliminate the specific latent features and mitigate the feature gap.
+Besides, we align features of adversarial samples in the fine-tuned model with features of natural samples in the naturally pre-trained model, further eliminating the specific latent features.
+Experiments demonstrate that our AFD alleviates the feature gap and improves the robustness.
+Contributions are summarized as follows:
+  • We observe the gap in features between natural and adversarial samples anomalously increases in adversarial fine-tuning methods, resulting from the specific latent features confused by adversarial perturbation.
+  • We provide the theoretical analysis of feature disentanglement.
+Then we introduce a disentanglement-based AFT approach, which explicitly models and further eliminates the specific latent features causing the feature gap by disentanglement and alignment.
   • Empirical evaluations show our approach mitigates the gap in features between natural and adversarial samples and surpasses existing methods.
 We also provide extended analyses for a holistic understanding.
-2 Related Work
+Related Work
+Adversarial Attack Adversarial Attack is first introduced in Goodfellow, Shlens, and Szegedy (2014), which fools DNNs by imperceptible perturbations.
+Madry et al.
+(2017) propose Projected Gradient Descent (PGD) attack based on the gradient projection and random start.
+Croce and Hein (2020) introduce a powerful adaptive attack (AutoAttack), including three white-box attacks and a black-box attack.
+Yu, Gao, and Xu (2021) demonstrate that exploiting latent features is highly effective against many defense techniques.
+Besides, adversarial attacks can also work in the physical world.
+Wei et al.
+(2022) present a comprehensive overview of physical adversarial attacks.
+Wei et al.
+(2023) propose a novel physical adversarial attack that applies the Warming Paste and Cooling Paste to hide persons from being detected.
+Liu et al.
+(2024) learn adversarial semantic perturbations in the latent space for high attack capabilities and low perceptibility.
+We take PGD and AutoAttack to evaluate robustness.
 Adversarial Training Adversarial training (AT) is the most effective technique to defend against attacks.
-Numerous studies have been conducted to boost robustness and improve the trade-off between accuracy and robustness.
 Madry et al.
-(2017) propose Projected Gradient Descent (PGD) attack and PGD-based adversarial training (PGD-AT), compelling the model to correctly classify adversarial samples within the epsilon sphere during training, which is the pioneer in adversarial learning.
+(2017) propose PGD-based adversarial training (PGD-AT), compelling the model to correctly classify adversarial samples within the epsilon sphere during training.
 Zhang et al.
 (2019) reduce the divergence of probability distributions of natural and adversarial samples to mitigate the difference between robust and natural accuracy.
+Xie et al.
+(2019) apply non-local means or other filters as the denoising block.
 Wang et al.
 (2020) find that misclassified samples harm adversarial robustness significantly, and propose to improve the model’s attention to misclassification by adaptive weights.
-Huang et al.
-(2020) replace labels with soft labels predicted by the model and adaptively reduce the weights of misclassified samples to alleviate the robust overfitting problem.
-Dong et al.
-(2021) also propose a similar idea of softening labels and explain the different effects of hard and soft labels on robustness by investigating the memory behavior of the model for random noisy labels.
+Yan et al.
+(2021) manipulates channels’ activations to align the channel with the relevance to predictions.
+Yang et al.
+(2021) introduce a disentanglement-based architecture to generate class-specific and class-irrelevant representations.
 Zhou et al.
 (2022) embed a label transition matrix into models to infer natural labels from adversarial noise.
-Though AT has good robust performances, it suffers from a large quantity of computing expenses and training costs.
-Some researchers have developed approaches to accelerate adversarial training, commonly known as fast adversarial training (FAT) (Wong et al., 2020; Kim et al., 2020; Huang et al., 2023).
+Kim et al.
+(2023) propose the FSR module to separate non-robust activations and recalibrate the features.
+Zhou et al.
+(2023) define two attributes of robust features to guide robust training.
+Though AT has good robust performances, it suffers from a large quantity of computing expenses.
+Some work has been developed to accelerate adversarial training, commonly known as fast adversarial training (FAT) (Wong, Rice, and Kolter 2020; Kim, Lee, and Lee 2020; Huang et al.
+2023).
 Since these FAT methods only require a little training time like AFT, we compare them with our method.
@@
 (2018) analyze the effect of AFT by comparing the decision boundary of a DNN before and after applying AFT.
-Some researchers (Tramer & Boneh, 2019; Maini et al., 2020; Madaan et al., 2021; Croce & Hein, 2022) aim to achieve multiple $l_{p}$ adversarial robustness through specific strategies.
-Jeddi et al.
-(2020) employ vanilla AFT to expedite the training of robust models with a few training epochs and a warm-up learning rate schedule.
+Kumari et al.
+(2019); Yu, Gao, and Xu (2021) discover the vulnerability of latent features in adversarial trained models and propose adversarial approaches to mitigate this.
+Jeddi, Shafiee, and Wong (2020) employ vanilla AFT to expedite the training of robust models with a few training epochs and a warm-up learning rate schedule.
+Some researchers (Tramer and Boneh 2019; Maini, Wong, and Kolter 2020; Madaan, Shin, and Hwang 2021; Croce and Hein 2022) aim to achieve multiple $l_{p}$ adversarial robustness through specific strategies.
 Based on representation learning, Suzuki et al.
-(2023) propose ARREST and AFT+RGKD, which reduce the feature gap between natural and adversarial samples.
-Zhu et al.
+(2023) constrain the representation to mitigate the accuracy-robustness trade-off.
+Although they aim to learn suitable features, the gap between natural and adversarial features can’t be eliminated by simple feature constraint, which has been indicated in Figure 1 (a)Zhu et al.
 (2023) propose a metric to measure the robustness of modules and fine-tune them for better out-of-distribution performance.
@@
 (2023) introduce a two-pipeline structure to improve the relationship between the weight norm and its gradient norm in batch normalization layers.
-Despite the extensive work, few methods (vanilla AFT, ARREST, and AFT+RGKD) are based on naturally pre-trained models, which are important baselines in our paper.
-3 Methodology
-In this section, we first reveal the issue of the increasing gap of features between natural and adversarial samples in AFT.
-Then we introduce a disentanglement-based method, to model and further remove the latent features leading to the feature gap.
-Our method is expected to mitigate the feature gap to enhance adversarial robustness.
-3.1 Preliminaries
+Despite the extensive work, few methods (vanilla AFT (Jeddi, Shafiee, and Wong 2020), ARREST and AFT+RGKD (Suzuki et al.
+2023)) are based on naturally pre-trained models, which are our important baselines.
+Figure 2: Overview of our AFD.
+Lock denotes the frozen parameters, $BZ$ denotes the batch size, $Cls$ denotes the class number, and numbers in brackets denote the sizes of features or parameters.
+During fine-tuning, Adversarial features are disentangled into $f_{1}$ and $f_{2}$.
+$\mathcal{L}_{CE}$ ensures $f_{2}$ to approximate confused features.
+We keep $f_{1}$ away from $f_{2}$ to eliminate confused features in $f_{1}$.
+Besides, we align $f_{1}$ with pre-trained intrinsic features $f_{nat}^{\prime}$ to further correct the prediction.
+During inference, we multiply $D_{1}$ and ${\omega}^{F}$ to obtain the robust model without additional modules.
+Preliminaries
 We denote a DNN as $h(x)=\omega\cdot g(x;\theta_{g})$, where $x$ is the input image, and $\theta_{g}$ is the parameters of the feature extractor $g$, i.e., the model before the last linear classifier.
@@
 The parameters of the naturally pre-trained and fine-tuned model are denoted as $\theta^{P}$ and $\theta^{F}$, respectively.
-The architecture and dataset for both pre-training and fine-tuning remain consistent, and the model parameters $\theta^{F}$ are initialized by $\theta^{P}$.
+The model parameters $\theta^{F}$ are initialized by $\theta^{P}$.
 $x$ denotes the clean sample, and $x^{\prime}$ denotes the adversarial sample.
-${\|{\delta}\|}_{p}\leq\epsilon$ denotes the adversarial perturbation, where $\epsilon$ represents the budget of adversarial attack.
-We use ${\|{\cdot}\|}_{p}$ to denote the $l_{p}$ norm.
+$\mathbb{N}(x,\epsilon)$ represents $\{x^{\prime}:\|x^{\prime}-x\|_{\infty}\leq\epsilon\}$, $\epsilon$ is the perturbation budget, ${\|{\cdot}\|}_{p}$ denotes the $l_{p}$ norm.
 Adversarial and natural features denote the features in the last hidden layer of adversarial and natural samples, i.e., $g(x^{\prime};\theta_{g})$ and $g(x;\theta_{g})$ respectively.
-$\mathcal{L}$ denotes the cross-entropy loss.
-3.2 Model the Latent Feature causing Feature Gap
-We evaluate the feature gap in AFT by measuring the ${L}_{p}$ distance between natural and adversarial features.
-As shown in Figure 2, the overall magnitudes of feature distance between natural and adversarial samples in three AFT methods (vanilla AFT, ARREST, AFT+RGKD) are becoming larger.
-The gap between the two types of features in ARREST initially decreases and then increases, which exhibits a V-shape trend.
-In addition, the feature gap in other AFT methods presents a consistently increasing trend.
-Despite AFT enhancing robust accuracy during training, the feature distance in AFT maintains a growing trend overall.
+Methodology
+In this section, we first reveal the issue of the increasing gap of features between natural and adversarial samples in AFT.
+Then we introduce a disentanglement-based method, to model and further remove the specific latent features leading to the feature gap.
+Our method is expected to mitigate the feature gap to enhance adversarial robustness.
+Specific Latent Feature Causing Feature Gap
+As shown in Figure 1 (a), there is a growing gap in features between natural and adversarial samples in existing AFT methods.
 Intuitively, the ideal robust model should have a consistent understanding and analysis of natural and adversarial samples, thus extracting similar (the same) features from them.
-Figure 2 indicates that existing AFT methods neglect the gap of features between natural and adversarial samples.
-Narrowing the feature gap is expected to achieve better performance and enhance robust accuracy.
+Thus, narrowing the feature gap is expected to achieve better robust performance.
 To realize the target, we can disentangle two types of features from the features of natural and adversarial samples.
-We define natural features $f_{nat}$ as the unconfused features $f_{u}$.
-We can define adversarial features $f_{adv}$ as the combination of the unconfused features $f_{u}$ and the confused features $f_{c}$, the latter of which leads to the gap in features between natural and adversarial samples.
-In this modeling, the unconfused features represent the features that contribute to accurate classification, which are easily extracted from natural samples.
-Besides, the confused features are identified as the latent features causing the gap between features of natural and adversarial samples, which are confused by the adversarial perturbation.
-As illustrated in Figure 1, if the confused features are removed in adversarial features, the features of adversarial samples can align with those of natural samples, resulting in a zero feature gap and accurate predictions for adversarial samples.
-Figure 2: $L_{p}$ distances of features between natural and adversarial samples on ResNet18 on CIFAR-10.
-The horizontal axis denotes the training epoch.
-Despite the initial decreasing fluctuations observed in ARREST and AFT+RGKD, AFT methods exhibit an increasing trend on the whole.
-We suppose the phenomenon in Figure 2 is attributable to the presence of the confused features $f_{c}$.
+We define natural features $f_{nat}$ as the intrinsic features $f_{i}$.
+We can define adversarial features $f_{adv}$ as the combination of the intrinsic features $f_{i}$ and the confused features $f_{c}$, the latter of which leads to the gap in features between natural and adversarial samples.
+In this modeling, the intrinsic features represent the features that contribute to accurate classification, which are easily extracted from natural samples.
+Besides, the confused features are identified as the specific latent features causing the gap between features of natural and adversarial samples, which are confused by the adversarial perturbation.
+As illustrated in Figure 1 (b), if the confused features are removed in adversarial features, the features of adversarial samples can align with those of natural samples, resulting in a zero feature gap and accurate predictions for adversarial samples.
+We suppose the phenomenon in Figure 1 (a) is attributable to the presence of the confused features $f_{c}$.
 Because the confused features persist as latent features and don’t diminish during fine-tuning, the gap between features of natural and adversarial samples cannot be bridged.
-The confused features induce the gap between natural and adversarial features, leading to the limited robust performance.
+The confused features induce the gap between natural and adversarial features, leading to limited robust performance.
 Based on our observations and analyses, we introduce a new approach named Adversarial Fine-tuning via Disentanglement (AFD), as shown in Figure 2.
 It applies a feature disentangler to explicitly extract the confused features, and remove them by the elimination and alignment of features.
-Figure 3: Overview of our AFD.
-Lock denotes the frozen parameters.
-The disentangler disentangles adversarial features into $f_{1}$ and $f_{2}$.
-We obtain confused features $f_{2}$ by matching the wrong predicted labels in $\mathcal{L}_{1}$.
-Next, we employ $\mathcal{L}_{2}$ to eliminate confused features from unconfused features $f_{1}$.
-Besides, we apply $\mathcal{L}_{3}$ to further align $f_{1}$ with pre-trained unconfused features $f_{nat}^{\prime}$ to correct the prediction.
-3.3 Adversarial Fine-tuning Based on Disentanglement
-The primary strategy is to explicitly disentangle the unconfused and confused features from the features of adversarial samples.
+Adversarial Fine-tuning Based on Disentanglement
+The primary strategy is to explicitly disentangle the intrinsic and confused features from the features of adversarial samples.
 We propose a feature disentangler to model confused features for further elimination.
 The features in the last hidden layer are input into the disentangler, yielding two disentangled feature vectors.
-We define these feature vectors as 1) $f_{1}$, features contributing to correct classification without confusion; and 2) $f_{2}$, features inducing wrong predictions of adversarial samples.
-To implement this disentangling, we introduce a disentangler with two linear blocks $D_{1},D_{2}$ between the feature extractor $g$ and the linear classifier $\omega$.
-These linear blocks are expected to disentangle adversarial features into the unconfused features and confused features.
-The disentangled feature vectors $f_{1}$ and $f_{2}$ are constrained to match the unconfused features and confused features, formulated as:
-$\begin{split}f_{1}&=\ {D_{1}}(g(x^{\prime};\theta_{g}^{F});\theta_{D_{1}}),\\ f_{2}&=\ {D_{2}}(g(x^{\prime};\theta_{g}^{F});\theta_{D_{2}}),\\ {\rm where}\ &\left\{\begin{array}[]{lr}f_{1}\stackrel{{\scriptstyle D_{1}}}{{\longrightarrow}}f_{u},&\\ f_{2}\stackrel{{\scriptstyle D_{2}}}{{\longrightarrow}}f_{c},&\end{array}\right.\end{split}$ (1)
-where $\theta_{g}^{F}$ denotes the parameters of the feature extractor in the fine-tuned model.
-We take $f_{1}$ to correctly classify samples and $f_{2}$ to match the confused features.
-The input and output dimensions of the linear block $D_{1},D_{2}$ are set to match the channel dimensions of the feature vector of adversarial samples before disentanglement.
+We define these feature vectors as intrinsic features $f_{1}$ and confused features $f_{2}$.
+To implement the disentanglement, we introduce a disentangler with two linear blocks $D_{1},D_{2}$ between the feature extractor $g$ and the linear classifier $\omega$.
+These linear blocks are expected to disentangle adversarial features into intrinsic and confused features.
+The disentangled feature vectors $f_{1}$ and $f_{2}$ are constrained to approximate the intrinsic and confused features.
+Given by:
+$\begin{split}&f_{1}={D_{1}}(g(x^{\prime};\theta_{g}^{F})),\\ &f_{2}={D_{2}}(g(x^{\prime};\theta_{g}^{F})),\\ &{\rm where}\ \left\{\begin{array}[]{lr}f_{1}\approx f_{i},&\\ f_{2}\approx f_{c},&\end{array}\right.\end{split}$ (1)
+where $\theta_{g}^{F}$ denotes the parameters of the feature extractor in the fine-tuned model. $x^{\prime}$ denotes the adversarial samples created by PGD-10 (Madry et al. 2017), formulated as:
+$x^{\prime}=\mathop{\boldsymbol{\Pi}}\limits_{\mathbb{N}(x,\epsilon)}\left(x+\epsilon\cdot\operatorname{sign}\left(\nabla_{x}\mathcal{L}_{CE}\left(\omega^{F}\cdot{D_{1}}(g(x;\theta_{g}^{F})),y\right)\right)\right),$ (2)
+where $\boldsymbol{\Pi}$ denotes a clamp function, $\operatorname{sign}$ denotes a sign function, $\mathcal{L}_{CE}$ denotes the cross-entropy loss.
+We take $f_{1}$ and $f_{2}$ to match the intrinsic and confused features.
+The input and output dimensions of $D_{1},D_{2}$ are set to match the channel dimensions of the feature vectors before disentanglement.
 This ensures the parameters of the linear block $\theta_{D_{1}}$ can be seamlessly integrated with the parameter of the last linear classifier $\omega^{F}$, expressed as $\theta_{D_{1}}\cdot\omega^{F}={\omega^{F\prime}}$.
-Thus we only need to load the linear classifier $\omega^{F\prime}$ without additional disentangling modules in the deployment (test) phase.
+Therefore, we only need to load the linear classifier $\omega^{F\prime}$ without additional disentangling modules in the inference (test) phase.
 To achieve good performance on both accuracy and robustness, we should design suitable loss functions to realize Formula 1.
-First, we suggest fitting $f_{2}$ to incorrectly predicted labels $y^{\prime}$ to model the confused features.
+We suggest fitting $f_{2}$ to incorrectly predicted labels $y^{\prime}$ to model the confused features.
 The confused features result in obvious shifts in both features and predicted probabilities, leading to wrong predicted class labels in most cases.
-Thus, it can be identified as the latent features causing the gap between natural and adversarial features.
-To disentangle the confused features from adversarial features, we force the disentangled features $f_{2}$ to match incorrect class labels, formulated as:
-$\begin{split}\mathcal{L}_{1}=\ &{\rm\mathop{arg\ min}\limits_{\theta_{D_{2}}}}\ \mathcal{L}(\omega^{F}\cdot f_{2},y^{\prime}),\\ {\rm where}\quad&y^{\prime}={\rm\mathop{arg\ sort}\limits_{y^{\prime}\neq y}}\ \omega^{F}\cdot f_{1},\end{split}$ (2)
-where $y^{\prime}$ denotes wrong predicted labels with the maximum predicted probability, and sort represents finding the index of the maximum value.
+Thus, it can be identified as the specific latent features causing the gap between natural and adversarial features.
+To decrease the dependence between the confused features and ground-truth labels, we force the disentangled features $f_{2}$ to match incorrect class labels, formulated as:
+$\begin{split}\mathcal{L}_{1}=\ &\mathcal{L}_{CE}(\omega^{F}\cdot f_{2},y^{\prime}),\\ {\rm where}\quad&y^{\prime}={\rm\mathop{max}\limits_{y^{\prime}\neq y}}\ (\omega^{F}\cdot f_{1}),\end{split}$ (3)
+where $y^{\prime}$ denotes wrong predicted labels with the maximum predicted probability.
 We only optimize the disentangling block $D_{2}$ instead of the entire model.
@@
 The exclusive force between $f_{1}$ and $f_{2}$ encourages the elimination of the confused features from the features of adversarial samples.
-This constraint aims to protect adversarial features from being perturbed by the confused features, and then gain the unconfused features $f_{1}$.
+This constraint aims to protect adversarial features from being perturbed by the confused features, and reduces the association between confused and intrinsic features.
 Given by:
-$\mathcal{L}_{2}={\rm\mathop{arg\ max}\limits_{\theta^{F},\theta_{D_{1}}}}\ \mathcal{D}(f_{1},\bar{f_{2}}),$ (3)
-where $\mathcal{D}$ denotes a distance function, $\bar{f_{2}}$ means the gradient back-propagation of $f_{2}$ is frozen.
-3.4 Adversarial Fine-tuning Based on Alignment
+$\mathcal{L}_{2}=\ -\mathcal{D}(f_{1},f^{-}_{2}),$ (4)
+where $\mathcal{D}$ denotes a distance function, $f^{-}_{2}$ means the gradient back-propagation of $f_{2}$ is frozen.
+Adversarial Fine-tuning Based on Alignment
 The second strategy is to align the adversarial features in the fine-tuned model with the natural features in the pre-trained model.
-Since the pre-trained model has a high natural accuracy and doesn’t suffer from adversarial attacks, the natural features in the pre-trained model can be identified as ideal unconfused features.
+Since the pre-trained model has a high natural accuracy and doesn’t suffer from adversarial attacks, the natural features in the pre-trained model can be identified as ideal intrinsic features.
 Thus, adversarial features in the fine-tuned model can be greatly improved by being aligned with the natural features in the pre-trained model.
-Besides, lots of work (Zhang et al., 2019; Wang et al., 2020; Suzuki et al., 2023) conducts an alignment between natural and adversarial features, which is regarded to mitigate the trade-off between accuracy and robustness.
+Besides, lots of work (Zhang et al.
+2019; Wang et al.
+2020; Suzuki et al.
+2023) conducts an alignment between natural and adversarial features, which is regarded to mitigate the trade-off between accuracy and robustness.
 We take the disentangled features $f_{1}$ to match the natural features in the pre-trained model, further removing the confused features from $f_{1}$.
 The AFT loss function with the alignment can be formulated as:
-$\mathcal{L}_{3}={\rm\mathop{arg\ min}\limits_{\theta^{F},\theta_{D_{1}}}}\ \mathcal{L}(\omega^{F}\cdot f_{1},y)+\gamma\cdot\mathcal{D}(f_{1},g(x;\theta_{g}^{P})),$ (4)
-where $\theta_{g}^{P}$ denotes parameters of the feature extractor in the pre-trained model, and $\gamma$ denotes the weight of the alignment.
+$\mathcal{L}_{3}=\mathcal{L}_{CE}(\omega^{F}\cdot f_{1},y)+\gamma\cdot\mathcal{D}(f_{1},g(x;\theta_{g}^{P})^{-}),$ (5)
+where $\theta_{g}^{P}$ denotes parameters of the feature extractor in the pre-trained model, $\gamma$ denotes the weight of the alignment, and $g(x;\theta_{g}^{P})^{-}$ denotes the gradient back-propagation of $g(x;\theta_{g}^{P})$ is frozen.
 Therefore, the objective function of our AFD is defined as follows:
 $\mathcal{L}_{total}=\alpha\cdot\mathcal{L}_{1}+\beta\cdot\mathcal{L}_{2}+\mathcal{L}_{3},$ (6)
-where $\alpha$ and $\beta$ denote the weights of $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$. The algorithm of AFD is as follows:
+where $\alpha$ and $\beta$ denote the weights of $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$. The algorithm of AFD in Algorithm 1.
 Algorithm 1 Adversarial Fine-tuning via Disentanglement
-0:  Training set $\mathcal{D}$, The $\mathcal{L}_{1}$ weight $\alpha$, The $\mathcal{L}_{2}$ weight $\beta$, Number of epochs $T$, Learning rate $\eta$, Inner maximization $\mathcal{A}$, Naturally pre-trained model parameters $\theta^{P}$, Model parameters except linear block $D_{2}$ parameters $\theta_{1}$, Linear block $D_{2}$ parameters $\theta_{2}$.
+0:  Training set $\mathcal{D}$, Hyperparameters $\alpha,\beta$, Number of epochs $T$, Learning rate $\eta$, Naturally pre-trained model parameters $\theta^{P}$, Model parameters except linear block $D_{2}$ parameters $\theta_{1}$, Linear block $D_{2}$ parameters $\theta_{2}$.
 0:  Fine-tuned model parameters $\theta_{1}$.
@@
      for $x,y\in\mathcal{D}$ do
-        $x^{\prime}=\mathcal{A}(x)$;
+        Calculate $x^{\prime}$ by by Formula 2;
         Calculate features $f_{1},f_{2}$ by Formula 1;
-        Calculate loss $\mathcal{L}_{1}$ by Formula 2;
+        Calculate loss $\mathcal{L}_{1}$ Formula 3;
         $\theta_{2}=\theta_{2}-\eta\cdot\frac{\mathrm{d}(\alpha\cdot\mathcal{L}_{1})}{\mathrm{d}\theta_{2}}$;
-        Calculate loss $\mathcal{L}_{2}$ by Formula 3;
+        Calculate loss $\mathcal{L}_{2}$ Formula 4;
         $\theta_{1}=\theta_{1}-\eta\cdot\frac{\mathrm{d}(\beta\cdot\mathcal{L}_{2})}{\mathrm{d}\theta_{1}}$;
-        Calculate loss $\mathcal{L}_{3}$ by Formula 4;
+        Calculate loss $\mathcal{L}_{3}$ by Formula 5;
         $\theta_{1}=\theta_{1}-\eta\cdot\frac{\mathrm{d}\mathcal{L}_{3}}{\mathrm{d}\theta_{1}}$;
@@
   end for
-4 Experiments
+Experiments
 This section presents experiments with AFD.
 We first introduce our experiment setting.
-We further make comparisons on multiple architectures and datasets to show our superiority.
+We further make comparisons on multiple architectures and datasets to show the effectiveness.
 Then we conduct an ablation study to show the functions of the disentangling module and alignment.
 Finally, we conduct an empirical analysis to support our hypothesis.
-4.1 Setting
-Datasets We conduct our experiments on three benchmark datasets including on CIFAR-10 and CIFAR-100 (Krizhevsky, 2009), Tiny-ImageNet (Deng et al., 2009).
+Setting
+Datasets We conduct experiments on three benchmark datasets including CIFAR-10 and CIFAR-100 (Krizhevsky 2009), Tiny-ImageNet (Deng et al.
+2009).
 CIFAR-10 dataset contains 60,000 color images having a size of $32\times 32$ in 10 classes, with 50,000 training and 10,000 test images.
@@
 Each class has 500 training images, 50 validation images and 50 test images.
-We only use training and validation images of Tiny-ImageNet in experiments.
-Baselines To make a comprehensive comparison, we have employed various techniques to demonstrate our superiority.
-Our primary baselines are AFT methods including vanilla AFT (vAFT) (Jeddi et al., 2020), AFT+RGKD (AFKD) (Suzuki et al., 2023), and ARREST (Suzuki et al., 2023).
+We only use training and validation images of Tiny-ImageNet.
+Baselines To make a comprehensive comparison, we have employed various techniques to demonstrate the effectiveness of our method AFD.
+Our primary baselines are AFT methods including vanilla AFT (vAFT) (Jeddi, Shafiee, and Wong 2020), AFT+RGKD (AFKD) (Suzuki et al.
+2023), and ARREST (Suzuki et al.
+2023).
 All of them are built upon naturally pre-trained models.
-Additionally, we compare our AFD with state-of-the-art AFT methods such as RiFT (Zhu et al., 2023) and TWINS (Liu et al., 2023), which rely on adversarially pre-trained models.
-Besides, we select several advanced fast adversarial training methods for comparison, including FreeAT (Shafahi et al., 2019), FGSM-GA (Andriushchenko & Flammarion, 2020), ATAS (Huang et al., 2023), since they are also low-cost techniques.
-Moreover, we compare with baseline methods of adversarial training including PGD-AT (Madry et al., 2017), TRADES (Zhang et al., 2019) to show our great advance in robustness.
-Optimization Details The optimizer used in all the experiments is Stochastic Gradient Descent (SGD) optimizer with a momentum of 0.9.
-We naturally pre-train all the models with a batch size of 128, training epochs of 100, and a weight decay of ${5.0\times 10^{-4}}$.
-The learning rate starts at 0.1 and then decays by $\times$0.1 with transition epochs $\{75,90\}$, following Zhang et al.
+Additionally, we compare our AFD with a state-of-the-art AFT method RiFT (Zhu et al.
+2023), which relies on adversarially pre-trained and has powerful out-of-distribution performances.
+Besides, we select several advanced fast adversarial training methods, including FreeAT (Shafahi et al.
+2019), FGSM-GA (Andriushchenko and Flammarion 2020), ATAS (Huang et al.
+2023), since they are also low-cost techniques.
+Moreover, we compare with baseline methods of adversarial training including PGD-AT (Madry et al.
+2017), TRADES (Zhang et al.
+2019) to show our great advance in robustness.
+Optimization Details The optimizer is Stochastic Gradient Descent (SGD) optimizer with a momentum of 0.9.
+We fine-tune the pre-trained models with a batch size of 128, a learning rate of 0.0025, a weight decay of ${5.0\times 10^{-4}}$, and training epochs of 20.
+The schedule of learning rate is the same as Suzuki et al.
 (2023).
-For AT methods, we train all the models with a batch size of 128, training epochs of 120, and a weight decay of ${2.0\times 10^{-4}}$.
-The learning rate starts at 0.1 and then decays by $\times$0.1 with transition epochs $\{75,90,100\}$, following Wang et al.
-(2020).
-For our AFD, we fine-tune the pre-trained models with a batch size of 128, an initial learning rate of 0.0025, a weight decay of ${5.0\times 10^{-4}}$, and training epochs of 20.
-The learning rate is changed to 0.0020 at the 11th epoch and is divided by half every two epochs thereafter, as Suzuki et al.
-(2023).
+We utilize Exponential Moving Average (EMA) (Hunter 1986) to gain better parameters.
 […]
-4.2 Main Results against White-Box Attack
-To evaluate AFD’s effectiveness, we conducted comparative experiments with other methods on CIFAR-10, CIFAR-100, and Tiny-ImageNet.
-We also make additional experiments on ResNet50 to fairly compare with TWINS.
-CIFAR-10 and CIFAR-100 As shown in Tables 1, methods above the double line denote AT and FAT methods, and methods under the double line denote AFT methods.
-It shows that among AFT methods, AFD exhibits the best robust performance and ARREST has the second best robust performance on ResNet18.
-Robust accuracies of AFD against AutoAttack (AA) surpass those of ARREST by 1.27% and 2.58% on CIFAR-10 and CIFAR-100, respectively.
-As indicated in Table 2, AFD exhibits the best robust performance and AFKT has the second best robust performance among all the methods on WideResNet-34-10.
-AFD achieves superior robust accuracies against AA compared to AFKD, with improvements of 2.05% and 2.34% on CIFAR-10 and CIFAR-100, respectively.
-In particular, RiFT doesn’t show great robust performances, since it is fine-tuned with natural data to enhance out-of-distribution performance.
-Besides, AFD with fewer training epochs outperforms AT methods.
-AFD surpasses PGD-AT in all respects.
-TRADES achieves a larger robust accuracy against AA on CIFAR-10 on ResNet18 than AFD by 1.02%.
-However, the situation is reversed on CIFAR-100 or on WideResNet-34-10, where AFD outperform TRADES against AA by 0.43%, 2.40%, and 2.37%.
-This can be attribute to the superiority of AFD and the robust over-fitting in AT.
-Moreover, FAT methods distinctly lag behind AFD in both natural and robust performances.
-Despite the advanced robust performance of AFD, its natural accuracy doesn’t surpass state-of-the-art AFT methods (ARREST).
-We make further discussion in the Section 4.3.
-In summary, AFD demonstrates superb robustness compared to AT, FAT and AFT methods.
-Table 1: Quantitative evaluations of AFD and other methods for CIFAR-10 and CIFAR-100 on ResNet18.
-Methods CIFAR-10 CIFAR-100
-Clean PGD AA Clean PGD AA
-PGD-AT 84.43 48.82 43.68 59.09 24.05 20.67
-TRADES 82.37 53.27 48.52 55.47 28.12 23.49
-${}^{{\dagger}}$ 78.37 40.90 36.00 50.56 19.57 15.09
-${}^{{\dagger}}$ 80.10 49.14 43.44 50.61 24.48 19.42
-${}^{{\dagger}}$ 81.22 50.03 45.38 55.49 27.68 22.62
-vAFT 83.93 51.40 45.48 60.83 26.55 21.96
-RiFT 84.49 49.24 43.66 59.51 24.01 20.48
-AFKD 84.71 51.32 46.09 65.54 28.09 22.92
-ARREST 85.71 51.23 46.23 67.00 26.89 21.34
-AFD 84.66 52.63 47.50 62.94 29.27 23.92
-Table 2: The performance of AFD and other methods on CIFAR-10 and CIFAR-100 on WideResNet-34-10.
-Methods CIFAR-10 CIFAR-100
-Clean PGD AA Clean PGD AA
-PGD-AT 86.35 49.57 45.81 60.30 25.14 22.65
-TRADES 85.68 52.53 49.22 57.67 28.26 25.25
-vAFT 87.42 53.57 48.49 64.92 27.19 23.59
-RiFT 84.49 49.24 45.83 60.59 25.57 22.67
-AFKD 88.54 54.86 49.57 68.50 30.12 25.28
-ARREST 88.42 54.16 49.45 69.52 29.47 24.55
-AFD 88.08 56.12 51.62 68.51 32.63 27.62
-Figure 4: The clean accuracy (NAT) and robust accuracy against PGD-20 (PGD) of AFD with different hyperparameters.
-Tiny-ImageNet We further conduct comparative experiments on a larger dataset, Tiny-ImageNet.
-As shown in Table 3, on ResNet18, AFD consistently outperforms all the methods in terms of robust performance and the robust performance of TRADES ranks second.
-The robust accuracy of AFD against PGD exceeds that of TRADES by 1.59% and its robust accuracy against AA outpaces that of TRADES by 1.58%.
-Moreover, on WideResNet-34-10, AFD keeps the best robust performance and AFKD exhibits the second best robust performance.
-The robust accuracy of AFD against PGD on WideResNet-34-10 surpasses that of AFKD by 2.56% and its robust accuracy against AA on WideResNet-34-10 is higher than that of AFKD by 2.76%.
-These findings indicate the superiority of AFD in terms of adversarial robustness on large datasets.
-Table 3: The performances of AFD and other methods on Tiny-ImageNet with ResNet18 and WideResNet-34-10 models.
-Methods ResNet18 WideResNet-34-10
-Clean PGD AA Clean PGD AA
-PGD-AT 52.12 19.38 16.79 54.24 19.03 16.51
-TRADES 49.28 22.59 17.06 49.22 23.33 18.51
-vAFT 52.75 21.29 16.76 57.17 22.24 18.02
-RiFT 52.24 20.89 17.05 55.09 19.51 16.79
-AFKD 57.17 22.29 16.68 60.92 24.27 19.01
-ARREST 60.36 19.28 12.58 65.07 22.73 16.67
-AFD 55.58 24.18 18.64 59.58 26.83 21.77
-ResNet50 We follow Liu et al.
-(2023) to make comparisons on ResNet50.
-Notice that TWINS employs a model adversarially pre-trained on ImageNet (Deng et al., 2009), and the number of its training periods is 60.
-We set the training epochs of AFD as 40.
-Results are shown in Table 4.
-Although TWINS benefits from prior knowledge of ImageNet and adversarial robustness of the pre-trained model, AFD still surpasses its robust performance on both CIFAR-10 and CIFAR-100.
-AFD achieves improvements of 1.54% and 1.56% under AA compared to TWINS, and its robust accuracies against PGD surpass those of TWINS by 1.88% and 2.52%.
-This demonstrates its superiority in terms of adversarial robustness on ResNet50.
-Table 4: The performances of AFD and the state-of-the-art method TWINS on CIFAR-10 and CIFAR-100 on ResNet50.
-Methods CIFAR-10 CIFAR-100
-Clean PGD AA Clean PGD AA
-${}^{{\dagger}}$ 89.77 52.24 48.46 69.48 28.52 23.47
-${}^{{\dagger}}$ 91.95 52.46 49.02 72.12 29.12 25.72
-AFD 86.31 54.34 50.56 65.48 31.64 27.28
-In summary, these experiments indicate that AFD always outperforms all the baselines for different adversarial attacks, datasets, and network architectures in the aspect of robustness.
-4.3 Ablation Study
-Effect of Disentangler As mentioned in Section 3.3, the disentangler should disentangles adversarial features into the confused and unconfused features, and further eliminate the confused features.
-As shown in Table 5, vAFT with the disentangler (vAFT+D) exhibits superior robustness compared to vAFT by 1.28% against PGD and by 1.40% against AA, respectively.
-It shows the feature disentangler effectively eliminates the confused features that lead to wrong predictions.
-Moreover, the natural accuracy of AFT-D outperforms that of AFT by 0.70%.
-This could be attributed to the disentangling module removing the class-irrelevant features of natural samples such as background and shadow.
-Effect of Alignment As shown in Table 5, vAFT with the alignment of features (vAFT+A) exhibits superior robustness compared to AFT against both PGD and AA by 0.39% and 0.72%, with a marginal increase in natural accuracy by 0.67%.
-It indicates that the alignment successfully assists to further remove the confused features.
-Effect of Combination The combination of disentangler and alignment (AFD) has the best performance on both natural and robust accuracy, which shows two strategies are compatible and the combination can boost the positive effects of individual strategies.
-However, in Table 1 and other tables, AFD exhibits a lower natural accuracy than state-of-the-art AFT methods.
-This is because the unconfused features $f_{1}$ have not been trained with natural samples, which may lead to an accuracy-robust trade-off.
+Methods CIFAR-10 CIFAR-100 Tiny-ImageNet
+Clean PGD AA Avg Clean PGD AA Avg Clean PGD AA Avg
+PGD-AT 84.43 48.82 43.68 58.98 59.09 24.05 20.67 34.60 52.12 19.38 16.79 29.43
+TRADES 82.37 53.27 48.52 61.39 55.47 28.12 23.49 35.69 49.28 22.59 17.06 29.64
+Free-AT 78.37 40.90 36.00 51.76 50.56 19.57 15.09 28.41 - - - -
+FGSM-GA 80.10 49.14 43.44 57.56 50.61 24.48 19.42 31.50 - - - -
+ATAS 81.22 50.03 45.38 58.88 55.49 27.68 22.62 35.26 - - - -
+vAFT 83.93 51.40 45.48 60.27 60.83 26.55 21.96 36.45 52.75 21.29 16.76 30.27
+RiFT 84.49 49.24 43.66 59.13 59.51 24.01 20.48 34.67 52.24 20.89 17.05 30.06
+AFKD 84.71 51.32 46.09 60.71 65.54 28.09 22.92 38.85 57.17 22.29 16.68 32.05
+ARREST 85.71 51.23 46.23 61.05 67.00 26.89 21.34 38.41 60.36 19.28 12.58 30.74
+AFD 84.88 52.70 47.40 61.66 65.69 28.56 23.05 39.10 57.41 23.63 17.73 32.92
+Table 1: Quantitative evaluations of all methods on ResNet18.
+Methods CIFAR-10 CIFAR-100 Tiny-ImageNet
+Clean PGD AA Avg Clean PGD AA Avg Clean PGD AA Avg
+PGD-AT 86.35 49.57 45.81 60.58 60.30 25.14 22.65 36.03 54.24 19.03 16.51 29.93
+TRADES 85.68 52.53 49.22 62.48 57.67 28.26 25.25 37.06 49.22 23.33 18.51 30.35
+vAFT 87.42 53.57 48.49 63.16 64.92 27.19 23.59 38.57 57.17 22.24 18.02 32.48
+RiFT 84.49 49.24 45.83 59.85 60.59 25.57 22.67 36.28 55.09 19.51 16.79 30.46
+AFKD 88.54 54.86 49.57 64.31 68.50 30.12 25.28 41.30 60.92 24.27 19.01 34.73
+ARREST 88.42 54.16 49.45 64.01 69.52 29.47 24.55 41.18 65.07 22.73 16.67 34.82
+AFD 88.08 56.12 51.62 65.27 68.51 32.63 27.62 42.92 63.89 25.73 19.13 36.25
+Table 2: Quantitative evaluations of all methods on WideResNet-34-10.
+Main Results
+ResNet18 First, AFD effectively mitigates the trade-off between generalization and robustness.
+As shown in Table 1, AFD takes the lead in robust performance and exhibits the second-best clean accuracy among AFT methods.
+It gains less clean accuracies compared to ARREST by 1.70%, but its robust accuracies against PGD-20 and AA outperform ARREST by 2.50% and 2.68%.
+The improvement in robustness is evidently larger than the drop in clean accuracy.
+This result indicates the positive effect of feature disentanglement on the accuracy-robustness trade-off.
+Moreover, AFD exceeds almost all the AT and FAT methods.
+Although the powerful AT baseline TRADES outperforms AFD in the robust accuracies on CIFAR-10 and CIFAR-100 by 0.56%, it surpasses TRADES in clean and average accuracy by 6.365% and 1.84%.
+Besides, it surpasses TRADES on Tiny-ImageNet in all metrics.
+The result implies that efficient AFD can achieve comparable performances with AT.
+WideResNet-34-10 Second, AFD takes further advantages on large architectures.
+As shown in Table 2, AFD maintains the best robust and average accuracies among all the approaches.
+It exceeds the second-best baselines by approximately 1.62% and 1.34%.
+The enhancement of robustness further enlarges compared to that on ResNet18.
+In addition, the natural generalization of AFD always ranks second and is less than the best clean accuracy by 0.88% on average.
+The discrepancy of clean accuracies between AFD and the best one is smaller than that on ResNet18 (1.69%).
+Furthermore, AFD outperforms AT baselines in all metrics.
+These demonstrate that AFD achieves a better trade-off between generalization and robustness on larger architectures.
+It’s worth noting that AFD doesn’t exhibit optimal natural generalization among AFT methods.
+It is attritued to merely adversarial features purified by disentangling modules.
+Intuitively, natural features can also be boosted by disentanglement mechanism.
+Nevertheless, it will result in additional computing cost.
+Besides, current AFD has remained the second-best clean accuracy among various methods.
+Summary These experiments indicate that AFD always outperforms all the baselines for different attacks, datasets, and architectures in the aspect of overall performance.
+It proves the effectiveness of our AFD.
+Ablation Study
+Disentangler and Alignment As shown in Table 3, vanilla AFT with the disentangler (vAFT+D) exhibits superior performances compared to vanilla AFT by 0.70%, 1.28%, and 1.40%, respectively.
+It shows that the disentangler effectively eliminates the confused features that lead to wrong predictions.
+Besides, vanilla AFT with the alignment of features (vAFT+A) outperforms vanilla AFT by 0.68%, 0.31%, and 1.33%.
+It indicates that the alignment successfully contributes to removing the confused features.
+Moreover, the combination of disentangler and alignment (AFD) has the best performance on both natural and robust accuracy, which shows two strategies are compatible and the combination can boost the positive effects of individual strategies.
+Methods vAFT vAFT+D vAFT+A AFD
+Clean 83.93 84.63 84.61 84.88
+PGD 51.40 52.68 51.71 52.70
+AA 45.48 46.88 46.81 47.40
 Table 3: The performances of vAFT and other methods on CIFAR-10 on ResNet18. ‘+D’ denotes the method with the disentangler, and ‘+A’ denotes the method with the alignment of features.
-Methods vAFT vAFT+D vAFT+A AFD
-Clean 83.93 84.63 84.61 84.66
-PGD 51.40 52.68 51.71 52.63
-AA 45.48 46.88 46.81 47.50
-Figure 5: $L_{p}$ distances of features between natural and adversarial samples. The horizontal axis represents the training epoch.
-Effect of Hyperparameter There are three hyperparameters in our AFD: $\alpha,\beta$ and $\gamma$.
-They control the weights of $\mathcal{L}_{1}$, $\mathcal{L}_{2}$, and the distance function in $\mathcal{L}_{3}$, respectively.
-We have conducted experiments with various combinations of weights on ResNet18 on CIFAR-10, as illustrated in Figure 4.
-There is a remarkable trade-off between accuracy and robustness when $\beta$ and $\gamma$ are growing.
-On the one hand, a larger value of $\beta$ indicates a larger exclusive force between the unconfused and confused features, leading to less non-robust features utilized in classification and better robustness.
+Hyperparameter There are three hyperparameters in our AFD: $\alpha,\beta$ and $\gamma$.
+As illustrated in Figure 4, there is a remarkable trade-off between accuracy and robustness when $\beta$ and $\gamma$ are growing.
+On the one hand, a larger value of $\beta$ indicates a larger exclusive force between the intrinsic and confused features, leading to less non-robust features utilized in classification and better robustness.
 On the other hand, as $\gamma$ increases, the natural accuracy increases but robust accuracy drops.
 This occurs because more feature information of natural samples is transferred to the fine-tuned model through the feature alignment.
-However, the variations of both natural and robust accuracy are less than 0.60%, demonstrating that our AFD is insensitive to variations of hyperparameters.
-4.4 Empirical Analysis of Disentangled Features
-We evaluate the performances of two types of disentangled features $f_{1}$, $f_{2}$ on various datasets and models.
-As shown in Table 6, the natural accuracies of $f_{2}$ are lower than those of $f_{1}$ by $4.53\%\sim 45.05\%$.
-It indicates the disentangler has successfully extracted and removed the confused features as $f_{2}$.
-High accuracies of $f_{1}$ indicates models benefit from eliminating the confused features.
-Interestingly, the disparity of accuracy on CIFAR-100 is smaller than that on CIFAR-10.
-We suppose that a dataset with more classes has more complex feature space, and it is more difficult to extract confused features by the linear disentangler.
-Maybe nonlinear disentanglers work better.
-Table 6: The natural accuracies of the unconfused features $f_{1}$ and confused features $f_{2}$.
-RN and WRN are abbreviations of ResNet and WideResNet, respectively.
-$\Delta$ denotes the accuracy disparity.
-Model Data $f_{1}$ $f_{2}$ $\Delta$
-RN18 CIFAR-10 84.66 39.61 45.05
-RN18 CIFAR-100 62.94 58.41 4.53
-RN50 CIFAR-10 83.18 47.35 35.83
-RN50 CIFAR-100 61.68 54.97 6.71
-WRN-34-10 CIFAR-10 88.08 46.89 41.19
-WRN-34-10 CIFAR-100 68.51 62.84 5.67
-Moreover, we analyze the feature distance between the unconfused features $f_{1}$ and natural features $f_{nat}$, as illustrated in Figure 5.
+Besides, the variations of both natural and robust accuracies are less than 0.60%, demonstrating that our AFD is insensitive to variations of hyperparameters.
+This conclusion is also true among different datasets.
+Figure 3: t-SNE visualization on CIFAR-10 on ResNet18.
+Figure 4: The clean accuracy (NAT) and robust accuracy against PGD-20 (PGD) of AFD with different hyperparameters.
+Empirical Analysis of Disentangled Features
+we analyze the feature distance between the intrinsic features $f_{1}$ and natural features $f_{nat}$, as shown in Figure 1 (a).
 In contrast to the growing trend of feature distance in existing AFT methods, the optimization trajectory of the feature distance in AFD exhibits a fast convergence trend.
@@
 Additionally, the magnitudes of the feature distance in AFD are smaller than other AFT methods, e.g., the $L_{\infty}$ distance of AFD (0.212) is notably less than those of ARREST and vanilla AFT (0.453, 0.236).
-These results demonstrates our disentanglement-based approach can enhance the adversarial robustness by alleviating the feature gap.
-5 Conclusion
+These results demonstrate our AFD can enhance the adversarial robustness by alleviating the feature gap.
+Moreover, we evaluate the performances of two types of disentangled features $f_{1}$, $f_{2}$ on various datasets and models.
+As shown in Table 4, the natural accuracies of $f_{2}$ are lower than those of $f_{1}$ by $5.67\%\sim 51.04\%$.
+It indicates the disentangler has successfully extracted and removed the confused features as $f_{2}$.
+High accuracies of $f_{1}$ indicate models benefit from eliminating the confused features.
+Interestingly, the disparity of accuracies on CIFAR-100 is smaller than those on CIFAR-10 and Tiny-ImageNet.
+We suppose that it is challenging to learn clear feature space from a small-resolution dataset with numerous classes, which made the linear disentangler difficult to extract confused features.
+We make further feature visualization by t-SNE (Van der Maaten and Hinton 2008).
+We take the fine-tuned model by AFD.
+The we acquire features of adversarial samples crafted by PGD-20.
+We visualizes two types of features: the features before disentanglement and the intrinsic features $f_{1}$.
+As illustrated in Figure 3, features after disentanglement has a distincter class boundary than features before disentanglement, indicating the effectiveness of disentanglement.
+Model Data $f_{1}$ $f_{2}$ $\Delta$
+RN18 CIFAR-10 84.88 39.05 45.83
+RN18 CIFAR-100 65.69 58.17 7.52
+RN18 Tiny-ImageNet 57.41 48.78 8.63
+RN50 CIFAR-10 87.01 35.97 51.04
+RN50 CIFAR-100 67.45 56.17 11.28
+WRN-34-10 CIFAR-10 88.08 46.89 41.19
+WRN-34-10 CIFAR-100 68.51 62.84 5.67
+WRN-34-10 Tiny-ImageNet 63.89 55.15 8.74
+Table 4: The natural accuracies of the intrinsic features $f_{1}$ and confused features $f_{2}$.
+RN and WRN are abbreviations of ResNet and WideResNet, respectively.
+$\Delta$ denotes the accuracy disparity.
+Training Time and Model Parameter
+[…]
+The vanilla ResNet18 and WideResNet-34-10 have parameters of 11.2M and 46.2M.
+In the training process, the additional module has a weight matrix that doesn’t depend on the input size.
+It is $2*512*512=0.5$M (4.46% increase) on ResNet18 and $2*1024*1024=2$M (4.33% increase) on WideResNet-34-10, which is acceptable for better training.
+Method Epoch ${\rm Time_{one}}$ ${\rm Time_{total}}$
+Standard 100 6.0 590
+PGD-AT 120 37 4440
+PGD-AT - - 4428
+TARDES 120 53 6360
+Free-AT 10 119 1188
+FGSM-GA 30 68 2052
+ATAS 30 36 1080
+vanilla AFT 20 37 1330
+RiFT 10 25 4690
+AFKD 20 40 1390
+ARREST 20 36 1310
+AFD 20 46 1510
+Table 5: Training time (second) of various methods on ResNet18 on CIFAR10.
+We show the time for one epoch (${\rm Time_{one}}$) and the expected values of the total time (${\rm Time_{total}}$).
+‘Standard’ denotes the natural pre-training.
+Limitation and futural work
+Subject to computation resources, we haven’t conducted experiments on ImageNet-1k.
+We will explore effective adversarial fine-tuning approaches on large-resolution datasets in the future.
+Besides, we observe that the increasing gap of features also appears in adversarial training.
+It remains a lot for further studies.
+Conclusion
 This paper uncovers a surprising increasing trend in the gap of features between natural and adversarial samples in AFT methods, and further investigates it from the perspective of features.
-We model features as two feature components, in which the confused features are defined as the latent features leading to the feature gap.
+We model features as the intrinsic features and confused features, in which the latter is defined as the specific latent features leading to the feature gap.
 Then we propose Adversarial Fine-tuning via Disentanglement (AFD) to bridge the feature gap to enhance robustness.
-Our approach contains two strategies: disentanglement and alignment.
-We design the feature disentangler to explicitly separate out the confused features from the features of adversarial samples, while the second strategy aligns the adversarial features with the natural features in the pre-trained model.
-Experiments demonstrate that AFD effectively mitigates the feature gap and achieves significant robust performance, even surpasses adversarial training methods.
-Future work will aim at the accuracy-robustness trade-off.
-Appendix A Performances of Pre-trained models
-Table 7: The natural accuracy of the naturally pre-trained models.
+We design the feature disentangler to explicitly separate out the confused features from features of adversarial samples.
+Besides, the disentangled features are aligned with the natural features in the pre-trained model.
+Experiments demonstrate that AFD effectively mitigates the feature gap and achieves a good trade-off between generalization and robustness.
+Acknowledgments
+This work was supported in part by the National Natural Science Foundation of China under Grants U22A2096, 62441601, and 62306227, in part by the Shaanxi Province Core Technology Research and Development Project under grant 2024QY2-GJHX-11, in part by the Fundamental Research Funds for the Central Universities under Grants QTZX23042 and ZYTS24142.
+Appendix A Further Discussion
+Question 1: What are the main differences of the motivation of this paper with the work to explore the existence of robust and non-robust features (Ilyas et al. 2019)?
+These papers have different assumptions and definitions.
+(1) For Ilyas et al.
+(2019), they assume the features of adversarial samples are composed of robust and non-robust features.
+Robust features facilitate robustness as well as generalization (i.e., natural accuracy), and non-robust features benefit generalization but damage robustness.
+(2) In our work, considering the inconsistency of features leads to the vulnerability of DNN, we assume the ideal robust model should have a consistent understanding of natural and adversarial samples, thus extracting similar features from them.
+Besides, we divide the features into intrinsic and confused features.
+Intrinsic features contain clean representation.
+Confused features are induced by adversarial noise and result in a feature gap between natural and adversarial examples, thus causing wrong predictions.
+The definitions of features are different from Ilyas et al.
+(2019).
+Question 2: As shown in Figure 1 (a), the vanilla AFT also has a low feature distance and its distance is less than the proposed method from epoch 2 to epoch 14.
+Does it mean that the vanilla AFT gains better robustness?
+It is because the vanilla AFT did not converge well in the early stages of training, allowing the corresponding adversarial noise to successfully attack with only a small feature distance.
+However, it doesn’t change the increasing trend of the feature gap.
+As shown in Figure 1 (a) after epoch 14, the feature distance of the vanilla AFT grows rapidly and exceeds AFD.
+Besides, its robust accuracy is lower than AFD all the time (an accuracy disparity of 1.2%-2.8%).
+When we enlarge the epochs to 40, our AFD rapidly converges to a small value in the feature distance and others still keep a low-speed increasing trend.
 Model Data Clean
@@
 WideResNet-34-10 Tiny-ImageNet 67.86
+Table 6: The clean accuracy of the pre-trained models.
+Appendix B Experiments of Pre-trained models
+We naturally pre-train all the models with a batch size of 128, training epochs of 100, and a weight decay of ${5.0\times 10^{-4}}$.
+The learning rate starts at 0.1 and then decays by $\times$0.1 with transition epochs $\{75,90\}$, following Zhang et al.
+(2019).
+For AT methods, we train all the models with a batch size of 128, training epochs of 120, and a weight decay of ${2.0\times 10^{-4}}$.
+The learning rate starts at 0.1 and then decays by $\times$0.1 with transition epochs $\{75,90,100\}$, following Wang et al.
+(2020).
 We conduct pre-training experiments on multiple datasets and architectures.
 As shown in Table 6, all of the pre-trained models have much higher accuracies than the fine-tuned models.
-It indicates that it is reasonable for feature alignment to regard features of the natural samples in the pre-trained models as the ideal unconfused features.
-Appendix B Feature Visualization
-Figure 6: t-SNE visualization of different methods. The proposed AFD shows a clearer separation in the area of the red box.
-To further demonstrate the effectiveness of AFD, Figure 6 visualizes the features of AFT methods using t-SNE (Van der Maaten & Hinton, 2008) on ResNet18 on CIFAR-10.
+It indicates that it is reasonable for feature alignment to regard features of the natural samples in the pre-trained models as the ideal instrinsic features.
+Figure 5: t-SNE visualization of different methods. Our AFD shows a clearer separation in the area of the red box.
+Appendix C Feature Visualization
+To further demonstrate the effectiveness of AFD, Figure 5 visualizes the features of AFT methods using t-SNE (Van der Maaten and Hinton 2008) on ResNet18 on CIFAR-10.
 We color each point to match its ground-truth label and highlight focused areas with red boxes.
@@
 This indicates that AFD successfully mitigates the confusion between features of distinct classes, leading to a more robust prediction.
-Appendix C Environments and Time of Experiments
-Table 8: The training time (second) of various methods on ResNet18 on CIFAR10. We show the time for one epoch (${\rm Time_{one}}$) and the expected values of the total time (${\rm Time_{total}}$).
-Method Epoch ${\rm Time_{one}}$ ${\rm Time_{total}}$
-Standard 100 6.0 590
-PGD-AT 120 37 4440
-${}^{{\dagger}}$ 4428
-TARDES 120 53 6360
-${}^{{\dagger}}$ 10 119 1188
-${}^{{\dagger}}$ 30 68 2052
-${}^{{\dagger}}$ 30 36 1080
-vanilla AFT 20 37 1330
-RiFT 10 25 4690
-AFKD 20 40 1390
-ARREST 20 36 1310
-AFD 20 46 1510
-[…]
-As shown in Table 8, TRADES costs the most time and ATAS costs the least time among all the methods.
-The time cost of RiFT surpasses those of other AFT methods because it needs an adversarially pre-trained model.
-Our AFD costs less time than TRADES by 4850 seconds and its total training time outperforms that of vanilla AFT by 180 seconds.
-Considering its advanced robustness, the time cost is worthy and acceptable.
+Appendix D Supplement
+We have conducted experiments on a machine with three NVIDIA GeForce RTX 4090 GPUs.
+We use a single GPU for ResNet models and three GPUs for WideResNet models.
+We take three runs for each fine-tuning experiment, and the deviation is 0.43%.
+The gap between features of natural and adversarial samples can be observed in all the hidden layers.
+We select features of the last hidden layer to illustrate the gap since the $l_{p}$ norm of the feature gap in the last hidden layer is the largest.
+We have evaluated the feature gap in AFT by measuring the ${L}_{1}$, ${L}_{2}$, and ${L}_{\infty}$ distances between natural and adversarial features.
+The trends among different metrics are similar to Figure 1 (a).
+The results in Figure 3 are obtained from the experiments on ResNet18 on CIFAR-10.
+The conclusions are similar among distinct datasets and models.
+The most important of the hyperparameters is $\gamma$, which can be searched in $\{10,25,30,35\}$.
+Since AFD is insensitive to variations of hyperparameters, it is convenient to apply AFD on various datasets and architectures without searching hyperparameters too many times.